In a golf tournament, there are twice as many amateur golfers as professional golfers. A total of 96 golfers are in the tournament. How many of them are amateurs?
Answer by nerdybill(7384)

In a golf tournament, there are twice as many amateur golfers as professional golfers. A total of 96 golfers are in the tournament. How many of them are amateurs?
.
Let A = number of amateur golfers
96-A = number of pro golfers
.
2A = 96-A
3A = 96
A = 96/3
A = 32 (amateurs)